### v2

@@ Abstract.
Efficient scene representations are essential for many computer graphics applications. A general unified representation that can handle both surfaces and volumes simultaneously, remains a research challenge. Inspired by recent methods for scene reconstruction that leverage mixtures of 3D Gaussians to model radiance fields, we formalize and generalize the modeling of scattering and emissive media using mixtures of simple kernel-based volumetric primitives. We introduce closed-form solutions for transmittance and free-flight distance sampling for different kernels, and propose several optimizations to use our method efficiently within any off-the-shelf volumetric path tracer. We demonstrate our method as a compact and efficient alternative to other forms of volume modeling for forward and inverse rendering of scattering media. Furthermore, we adapt and showcase our method in radiance field optimization and rendering, providing additional flexibility compared to current state of the art given its ray-tracing formulation. We also introduce the Epanechnikov kernel and demonstrate its potential as an efficient alternative to the traditionally-used Gaussian kernel in scene reconstruction tasks. The versatility and physically-based nature of our approach allows us to go beyond radiance fields and bring to kernel-based modeling and rendering any path-tracing enabled functionality such as scattering, relighting and complex camera models.
Volumetric Primitives, Scattering, Radiance Fields
submissionid: 500copyright: none
@@ 1. Introduction
Volumetric representations of appearance have found good success at representing complex appearances such as cloth and hair (Schröder et al., 2011; Zhao et al., 2011; Khungurn et al., 2015; Aliaga et al., 2017), trees (Neyret, 1998; Loubet and Neyret, 2017), clouds and smoke (Kallweit et al., 2017), or particle aggregates (Moon et al., 2007; Meng et al., 2015; Müller et al., 2016), to name a few. Most of these approaches leverage grid-like representations of appearance such as voxels-grids, potentially with some multilevel underlying hierarchy, for which highly efficiently libraries exist (DreamWorksAnimation, 2023). Voxel grids are standard in rendering, and used in most research (Jakob, 2010; Jakob et al., 2022b; Pharr et al., 2023) and production renderers (Georgiev et al., 2018; Christensen et al., 2018; Burley et al., 2018; Fascione et al., 2018).
@@
Voxel-grids are flexible, provide $O(1)$ query time, and spatial connectivity for e.g., downsampling and level-of-detail representations (Heitz et al., 2015; Loubet and Neyret, 2017). Unfortunately, they scale poorly in terms of memory, specially when representing very fine detail and sparse structures. While hierarchical grid-based methods (Fong et al., 2017) have proven very effective at subdividing complex heterogeneous models into simpler local components, their building blocks are still grid-based, essentially inheriting the same problems. For an efficient integration of e.g., transmittance, voxel grids require the use of stochastic tracking techniques (Novák et al., 2018), which introduces additional variance when rendering media.
@@
In parallel, volume rendering techniques and continuous volumetric representations have recently seen unprecedented interest in the fields of computer vision and image-based graphics, spearheaded by efforts such as neural radiance fields (Mildenhall et al., 2020). These methods allow to capture and render photorealistic three-dimensional scenes, by optimizing an underlying volumetric representation of matter. Different volumetric representations have been proposed, including point-based methods (Kerbl et al., 2023; Sainz and Pajarola, 2004), implicit neural models (Mildenhall et al., 2020; Barron et al., 2021, 2022), explicit voxel-based representations (Fridovich-Keil et al., 2022; Yu et al., 2021), and hybrid approaches (Müller et al., 2022; Xu et al., 2022; Lombardi et al., 2021). While implicit neural models generally achieve the highest compression rates, point-based methods can be considered as the state of the art in terms of image quality and rendering speed for radiance fields rendering, while still being substantially more compact than voxel-grid structures. However, while their derivation starts from physics (i.e., the volume rendering equation), these methods in general are not suitable for their use within a physics-based renderer, given the fairly large assumptions and simplifications imposed in their image formation model to favour speed and ease of optimization.
@@
Inspired by the success of recent 3D Gaussian-based representations of radiance fields (Kerbl et al., 2023), we propose a new representation for general scattering and emissive media based on mixtures of three-dimensional kernel-based volumetric primitives. Each primitive statistically represents a spatial distribution of matter with the same optical properties. We introduce this representation inside the radiative transfer framework (Chandrasekhar, 1960) and derive closed-form solutions for transmittance and emission without the need of stochastic sampling, as well as sampling routines for computing inscattering.
While the scene representation is similar to the Gaussian splatting technique (Kerbl et al., 2023), the image formation model is fundamentally different: Kerbl et al. assume that each primitive becomes a billboard oriented towards the camera and therefore the density along the ray becomes a sum of delta functions; this allows them to implement their model in an extremely efficient rasterization pipeline. In contrast, our radiative primitive-based formulation preserves the primitive’s three-dimensional density and integrates following the physical transport process. This is crucial for physics-based rendering since it allows preserving reciprocity, while solving some view-dependent primitive ordering problems in Gaussian splatting.
@@
Our theoretical framework is general and supports a wide range of three-dimensional kernels. We demonstrate our work using the Gaussian kernel employed by Kerbl et al. (2023), but also on 3D Epanechnikov kernels, implementing closed-form solutions for volumetric transmittance, emission, and free-flight sampling. We implement our work by leveraging hardware-accelerated ray tracing for finding the primitives’ bounds and integrating their contribution along the ray. This fits well inside modern ray-tracing-based renderers, allowing us to compute transmittance and emission for both primary and secondary rays, needed for multiple scattering. We also derive the adjoint of our image formation model, which we demonstrate in inverse rendering applications. We showcase our primitive-based volumetric model through several applications both in forward and inverse rendering: 1) traditional forward rendering of scattering media, where we demonstrate improved performance over voxel-grids using local aggregated statistics and state of the art transmittance estimators; 2) inverse rendering of scattering and purely absorptive media; 3) inverse tomographic reconstructions from focus stacks using telecentric cameras and 4) radiance field optimization and rendering of complex scenes (both real and synthetic), with increased generality and controllability. Furthermore, our integration into the general RTE allows for future extensions in relighting and scattering. In short, our contributions are:
@@ 2. Related Work
Simulating light transport in media has been thoroughly studied in computer graphics (Novák et al., 2018). It involves solving the radiative transfer equation (RTE) (Chandrasekhar, 1960), which has been extended to account for anisotropic (Jakob et al., 2010), refractive (Ament et al., 2014), or spatially correlated (Jarabo et al., 2018; Bitterli et al., 2018) media. These generalizations make it possible to render a wider range of light transport phenomena.
Numerical estimation of transmittance is at the core of modern volumetric light transport simulation (Novák et al., 2018); via biased quadrature rules (ray-marching) (Tuy and Tuy, 1984; Muñoz, 2014) and modern unbiased variants (Kettunen et al., 2021), to stochastic tracking algorithms (null-scattering estimators) (Woodcock, 1965), including extensions with variance reduction via control variates (Novák et al., 2014; Szirmay-Kalos et al., 2017; Crespo et al., 2021), alternative formulations based on power-series (Georgiev et al., 2019), or differentiable formulations targeting inverse rendering (Nimier-David et al., 2022). All these approaches require tight estimates of the maximum density (majorant) for performance, either obtained in precomputation (Yue et al., 2010; Szirmay-Kalos et al., 2011) or estimated on-the-fly (Carter et al., 1972; Kutz et al., 2017; Galtier et al., 2013; Misso et al., 2023). Our approach, on the other hand, allows to compute transmittance in closed-form, as the product of transmittance of all primitives along the ray.
@@
For heterogeneous media, discrete voxel-grid approaches are ubiquitously found in most applications due to their simplicity and flexibility. They can feature multi-level hierarchical structures (Museth, 2013, 2021), which are helpful for faster traversal and filtering for level-of-detail applications, but suffer from poor scalability in terms of memory consumption when the modeled medium possesses finer details. Sparser representations like collections of isotropic and simple volumes have been used in other fields, e.g., particle physics (Brown and Martin, 2003), where they can model neutron transport in graphite pebble-bed reactors. The advantage of having collections of these simple volumes is that, individually, they offer closed-form solutions for transmittance estimation and sampling, reducing variance when compared with most tracking methods normally used in voxel grids. However, searching for the boundaries of these volumes can be inefficient and slow in many situations (Bitterli et al., 2018). In graphics, we can find hierarchical grid-based approaches (Fong et al., 2017) for subdividing complex, potentially sparse, volumes into smaller, more compact representations, though eventually they rely on smaller voxel-based primitives. Alternatively, implicit representations of participating media have been proposed, most notably in the context of radiance fields but also in compression of large volumes, including large multilayer perceptrons (MLP) (Mildenhall et al., 2020; Kim et al., 2024) or sparse hash grids of features combined with tiny MLPs (Müller et al., 2022). These approaches result in large compression rates, at the cost of expensive queries. Closer to our work, Knoll et al. (2021) proposed to model emissive and absorbing media using isotropic Gaussian mixture models, in the context of particle-based volumes (Max, 1979); however, they did not leverage closed-forms expressions for transmittance, requiring stochastic integration via tracking, and cannot represent scattering media. In contrast, our work proposes closed-form solutions for transmittance and sampling, and further extends to support scattering media, anisotropic kernels, and efficient inverse rendering.
@@
Early works on inverse rendering of heterogeneous media used inverse volumetric rendering for matching micron-scale cloth patches to measures (Khungurn et al., 2015), approximating multiscale volumetric representations (Zhao et al., 2016), scattering compensation for 3D printing (Elek et al., 2017; Sumin et al., 2019; Nindel et al., 2021; Condor et al., 2023), or inverse scattering (Gkioulekas, 2016) by using hand-made derivatives of light transport. Zhang et al. (2019) proposed a differential radiative transfer framework suitable for inverse volumetric rendering. These works use voxel grids for representing media density, though our proposed representation would fit in these inverse pipelines.
@@
Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) introduced a continuous, implicit volumetric representation based on multi-layer perceptrons (MLPs). Subsequent works have leveraged multi-scale representations (Barron et al., 2021) and extended to unbounded scenes (Barron et al., 2022; Zhang et al., 2020; Reiser et al., 2023); however, training and evaluating the MLPs still requires significant computation. In contrast, Instant-NGP (Müller et al., 2022) relies on a sparse hash grid of features and a small MLP, enabling faster scene reconstruction. Sparsification and hierarchization has been a popular avenue for optimizing radiance field fitting and rendering (Yu et al., 2021; Hedman et al., 2021; Božič et al., 2022; Chen et al., 2022; Lombardi et al., 2021), achieving real-time rendering even on mobile devices. Most recently, 3D Gaussian Splatting (Kerbl et al., 2023) removes the need for MLPs altogether, relying instead on an explicit Gaussian-based representation. Combined with an efficient splatting technique for rasterization (Zwicker et al., 2001), their approach is able to achieve fast rendering performance and high-quality reconstruction, allocating more primitives in areas where finer detail is needed. Our work takes Kerbl’s approach, and extends it to general scattering media in the context of the radiative transfer framework.
@@
Gaussian mixture models defined over the sphere have been extensively used in graphics for fitting BRDFs or environment maps (Xu et al., 2013), precomputed radiance transfer (Green et al., 2006), or on-line path guiding (Vorba et al., 2014). Yan et al. (2016) fitted a 4D Gaussian mixture model for accelerating computation when rendering glinty surfaces. Jakob et al. (2011) proposed to fit a 3D anisotropic Gaussian mixture model to the distribution of radiance in media, showing accelerated rendering using photon-based techniques. Despite their work sharing some similarities to ours (closed-form integration along Gaussians), it focuses on fitting radiance in the context of radiance estimation, while our approach focuses on modeling matter (density) and can be used in arbitrary volumetric renderers. Combining both representations (kernel-based representations for both media and radiance) would be an interesting line of future work.
@@ 3. Background
Under the geometric optics assumption, light transport in participating media is governed by the radiative transfer equation (RTE) (Chandrasekhar, 1960). We compute the incident radiance $L(\mathbf{x},\omega)$ at point $\mathbf{x}$ and direction $\omega$ as
@@
that models the light scattered in direction $\omega$ as a function of the integral of the light incoming from all directions on the unit sphere ${\mathcal{S}^{2}}$. The interaction of light and matter is characterized by the optical properties of the medium, described in terms of the differential probabilities of absorption and scattering (or coefficients) $\mu_{a}(\mathbf{x})$ and $\mu_{s}(\mathbf{x})$ respectively, the extinction coefficient $\mu_{t}(\mathbf{x})=\mu_{a}(\mathbf{x})+\mu_{s}(\mathbf{x})$, the phase function $f_{\text{p}}(\mathbf{x},\omega^{\prime}\rightarrow\omega)$, and the emission term $Q(\mathbf{x},\omega)$. Assuming uncorrelated statistics of matter, the differential probability of extinction is defined as the product between the density $\rho(\mathbf{x})$ of the particles forming the medium at point $\mathbf{x}$ [m-3] and the particles’ cross section $\sigma(\mathbf{x})$ [m2] as $\mu_{t}(\mathbf{x})=\rho(\mathbf{x})\,\sigma(\mathbf{x})$. For simplicity, we ignore the wavelength dependency in Equation (1), assume scalar elastic light-matter interactions, and omit the potential anisotropy of the medium (Jakob et al., 2010); generalizing the following derivations to include these additional dimensions is trivial.
By integrating both sides of the differential RTE (1) along the direction $\omega$ we obtain the volumetric rendering equation, that models radiance $L(\mathbf{x},\omega)$ as
$L(\mathbf{x}_{0},\omega)$ $=\int_{0}^{s}\textrm{T}(\mathbf{x},\mathbf{x}_{t})\,\mu_{t}(\mathbf{x}_{t})\,L_{o}(\mathbf{x}_{t},\omega)\,\text{d}t,$
(3) $+\textrm{T}(\mathbf{x},\mathbf{x}_{s})L(\mathbf{x}_{s},\omega)\qquad\text{ with }$
@@
Finally, $L(\mathbf{x}_{s},\omega)$ is the outgoing radiance at the medium boundary point $\mathbf{x}_{s}$ at distance $s$, which is modeled by the rendering equation (Kajiya, 1986).

### v3

@@ Abstract.
Efficient scene representations are essential for many computer graphics applications. A general unified representation that can handle both surfaces and volumes simultaneously, remains a research challenge. In this work we propose a compact and efficient alternative to existing volumetric representations for rendering such as voxel grids. Inspired by recent methods for scene reconstruction that leverage mixtures of 3D Gaussians to model radiance fields, we formalize and generalize the modeling of scattering and emissive media using mixtures of simple kernel-based volumetric primitives. We introduce closed-form solutions for transmittance and free-flight distance sampling for different kernels, and propose several optimizations to use our method efficiently within any off-the-shelf volumetric path tracer. We demonstrate our method in both forward and inverse rendering of complex scattering media. Furthermore, we adapt and showcase our method in radiance field optimization and rendering, providing additional flexibility compared to current state of the art given its ray-tracing formulation. We also introduce the Epanechnikov kernel and demonstrate its potential as an efficient alternative to the traditionally-used Gaussian kernel in scene reconstruction tasks. The versatility and physically-based nature of our approach allows us to go beyond radiance fields and bring to kernel-based modeling and rendering any path-tracing enabled functionality such as scattering, relighting and complex camera models. Code is available at https://github.com/facebookresearch/volumetric_primitives
Volume Rendering, Scattering, Radiance Fields, 3D Reconstruction, Volumetric Primitives, Volumetric Representations, Ray Tracing, Inverse Rendering
copyright: ccjournal: TOGjournalyear: 2025journalvolume: 44journalnumber: 1article: 10publicationmonth: 2doi: 10.1145/3711853ccs: Computing methodologies Shape representationsccs: Computing methodologies Ray tracingccs: Computing methodologies Volumetric modelsccs: Computing methodologies Point-based modelsccs: Computing methodologies Reconstruction
@@ 1. Introduction
Volumetric representations of appearance have found good success at representing complex appearances such as cloth and hair (shroder11; zhao11; khungurn2015; aliaga17), trees (neyret98; loubet2017), clouds and smoke (kallweit2017), or particle aggregates (moon07; meng15; mueller16efficient), to name a few. In this work we propose a compact volumetric representation of appearance, that allows efficient physics-based rendering of both scattering and emissive media.
@@
Most previous approaches for modeling volumetric appearances leverage grid-like representations of appearance such as voxels-grids, potentially with a multilevel underlying hierarchy, for which highly efficiently libraries exist (museth2013vdb). Voxel grids are standard in rendering, and used in most research (jakob2010mitsuba; jakob2022mitsuba3; pbrt) and production renderers (fajardo18arnold; renderman; hyperion; manuka). Voxel-grids are flexible, provide $O(1)$ query time, and spatial connectivity for e.g., downsampling and level-of-detail representations (heitz2015sggx; loubet2017). Unfortunately, they scale poorly in terms of memory, specially when representing very fine detail and sparse structures. While hierarchical grid-based methods (ProductionVolumeRendering2017) have proven very effective at subdividing complex heterogeneous models into simpler local components, their building blocks are still grid-based, essentially inheriting the same problems. For an efficient integration of e.g., transmittance, voxel grids require the use of stochastic tracking techniques (novak2018monte), which introduces additional variance when rendering media.
@@
In parallel, volume rendering techniques and continuous volumetric representations have recently seen unprecedented interest in the fields of computer vision and image-based graphics, spearheaded by efforts such as neural radiance fields (mildenhall2020). These methods allow to capture and render photorealistic three-dimensional scenes, by optimizing an underlying volumetric representation of matter. Different volumetric representations have been proposed, including point-based methods (kerbl20233Dgaussians; sainz04), implicit neural models (mildenhall2020; mipnerf; barron2022mipnerf360), explicit voxel-based representations (yu21plenoxels; yu2021plenoctrees), and hybrid approaches (mueller2022instant; pointnerf; Lombardi21). While implicit neural models generally achieve the highest compression rates, point-based methods can be considered as the state of the art in terms of image quality and rendering speed for radiance fields rendering, while still being substantially more compact than voxel-grid structures. However, while their derivation starts from physics (i.e., the volume rendering equation), these methods in general are not suitable for their use within a physics-based renderer, given the fairly large assumptions and simplifications imposed in their image formation model to favour speed and ease of optimization.
@@
Inspired by the success of recent 3D Gaussian-based representations of radiance fields (kerbl20233Dgaussians), we propose a new representation for general scattering and emissive media based on mixtures of three-dimensional kernel-based volumetric primitives. Each primitive statistically represents a spatial distribution of matter with the same optical properties. We introduce this representation inside the radiative transfer theory (RTT) (chandrasekhar1960radiative) and derive closed-form solutions for transmittance and emission without the need of stochastic sampling, as well as sampling routines for computing inscattering.
@@
While the scene representation is similar to the Gaussian splatting technique (kerbl20233Dgaussians), the image formation model is fundamentally different: Kerbl et al. assume that each primitive becomes a billboard oriented towards the camera and therefore the density along the ray becomes a sum of delta functions; this allows them to implement their model in an extremely efficient rasterization pipeline. In contrast, our radiative primitive-based formulation preserves the primitive’s three-dimensional density and integrates following the physical transport process. This is crucial for physics-based rendering since it allows preserving reciprocity, while solving some view-dependent primitive ordering problems in Gaussian splatting. Our theoretical framework is general and supports a wide range of three-dimensional kernels. We demonstrate our work using the Gaussian kernel employed by Kerbl et al. (kerbl20233Dgaussians), but also on 3D Epanechnikov kernels, implementing closed-form solutions for volumetric transmittance, emission, and free-flight sampling. We implement our work by leveraging hardware-accelerated ray tracing for finding the primitives’ bounds and integrating their contribution along the ray. This fits well inside modern ray-tracing-based renderers, allowing us to compute transmittance and emission for both primary and secondary rays, needed for multiple scattering. We also derive the adjoint of our image formation model, which we demonstrate in inverse rendering applications. We showcase our primitive-based volumetric model through several applications both in forward and inverse rendering: 1) traditional forward rendering of scattering media, where we demonstrate improved performance over voxel-grids using local aggregated statistics and state of the art transmittance estimators; 2) inverse rendering of scattering and purely absorptive media; 3) inverse tomographic reconstructions from focus stacks using telecentric cameras and 4) radiance field optimization and rendering of complex scenes (both real and synthetic), with increased generality and controllability. Furthermore, our integration into the general RTT allows for future extensions in relighting and scattering. In short, our contributions are:
@@ 2. Related Work
Simulating light transport in media has been thoroughly studied in computer graphics (novak2018monte). It involves solving the radiative transfer equation (RTE) (chandrasekhar1960radiative), which has been extended to account for anisotropic (jakob2010radiative), refractive (ament14refractive), or spatially correlated (jarabo2018radiative; bitterli2018radiative) media. These generalizations make it possible to render a wider range of light transport phenomena.
Numerical estimation of transmittance is at the core of modern volumetric light transport simulation (novak2018monte); via biased quadrature rules (ray-marching) (tuy1984direct; munoz2014higher) and modern unbiased variants (kettunen21), to stochastic tracking algorithms (null-scattering estimators) (woodcock), including extensions with variance reduction via control variates (novak14; szirmay2017unbiased; crespo2021primary), alternative formulations based on power-series (georgiev19), or differentiable formulations targeting inverse rendering (nimierdavid2022unbiased). All these approaches require tight estimates of the maximum density (majorant) for performance, either obtained in precomputation (yue2010unbiased; szirmay2011free) or estimated on-the-fly (carter1972monte; kutz2017spectral; galtier2013integral; misso2023progressive). Our approach, on the other hand, allows to compute transmittance in closed-form, as the product of transmittance of all primitives along the ray.
@@
For heterogeneous media, discrete voxel-grid approaches are ubiquitously found in most applications due to their simplicity and flexibility. They can feature multi-level hierarchical structures (museth2013vdb; museth2021nanovdb), which are helpful for faster traversal and filtering for level-of-detail applications, but suffer from poor scalability in terms of memory consumption when the modeled medium possesses finer details. Sparser representations like collections of isotropic and simple volumes have been used in other fields, e.g., particle physics (Brown03), where they can model neutron transport in graphite pebble-bed reactors. The advantage of having collections of these simple volumes is that, individually, they offer closed-form solutions for transmittance estimation and sampling, reducing variance when compared with most tracking methods normally used in voxel grids. However, searching for the boundaries of these volumes can be inefficient and slow in many situations (bitterli2018radiative). In graphics, we can find hierarchical grid-based approaches (ProductionVolumeRendering2017) for subdividing complex, potentially sparse, volumes into smaller, more compact representations, though eventually they rely on smaller voxel-based primitives. Alternatively, implicit representations of participating media have been proposed, most notably in the context of radiance fields but also in compression of large volumes, including large multilayer perceptrons (MLP) (mildenhall2020; neuralvdb) or sparse hash grids of features combined with tiny MLPs (mueller2022instant). These approaches result in large compression rates, at the cost of expensive queries. Closer to our work, Knoll et al. (Knoll2021) proposed to model emissive and absorbing media using isotropic Gaussian mixture models, in the context of particle-based volumes (max1979atomlll); however, they did not leverage closed-forms expressions for transmittance, requiring stochastic integration via tracking, and cannot represent scattering media. In contrast, our work proposes closed-form solutions for transmittance and sampling, and further extends to support scattering media, anisotropic kernels, and efficient inverse rendering.
@@
Early works on inverse rendering of heterogeneous media used inverse volumetric rendering for matching micron-scale cloth patches to measures (khungurn2015), approximating multiscale volumetric representations (zhao2016downsample), scattering compensation for 3D printing (elek17; sumin19; nindel2021gradient; Condor2023), or inverse scattering (Gkioulekas2016AFF) by using hand-made derivatives of light transport. Zhang et al. (Zhang2019drt) proposed a differential radiative transfer framework suitable for inverse volumetric rendering. These works use voxel grids for representing media density, though our proposed representation would fit in these inverse pipelines.
@@
Neural Radiance Fields (NeRF) (mildenhall2020) introduced a continuous, implicit volumetric representation based on multi-layer perceptrons (MLPs). Subsequent works have leveraged multi-scale representations (mipnerf) and extended to unbounded scenes (barron2022mipnerf360; zhang2020nerf++; reiser2023merf); however, training and evaluating the MLPs still requires significant computation. In contrast, Instant-NGP (mueller2022instant) relies on a sparse hash grid of features and a small MLP, enabling faster scene reconstruction. Sparsification and hierarchization has been a popular avenue for optimizing radiance field fitting and rendering (yu2021plenoctrees; hedman2021snerg; bovzivc2022neu=ral; chen2022mobilenerf; Lombardi21), achieving real-time rendering even on mobile devices, or helping on accelerating off-line rendering with complex assets (zhu2021neural; condor22). Most recently, 3D Gaussian Splatting (kerbl20233Dgaussians) removes the need for MLPs altogether, relying instead on an explicit Gaussian-based representation. Combined with an efficient splatting technique for rasterization (ewasplattingzwicker04), their approach is able to achieve fast rendering performance and high-quality reconstruction, allocating more primitives in areas where finer detail is needed. Our work takes Kerbl’s approach, and extends it to general scattering media in the context of the radiative transfer framework.
@@
Gaussian mixture models defined over the sphere have been extensively used in graphics for fitting BRDFs or environment maps (Xu13sigasia), precomputed radiance transfer (green2006view), or on-line path guiding (vorba2014line). Yan et al. (yan2016position) fitted a 4D Gaussian mixture model for accelerating computation when rendering glinty surfaces. Jakob et al. (jakob2011progressive) proposed to fit a 3D anisotropic Gaussian mixture model to the distribution of radiance in media, showing accelerated rendering using photon-based techniques. Despite their work sharing some similarities to ours (closed-form integration along Gaussians), it focuses on fitting radiance in the context of radiance estimation, while our approach focuses on modeling matter (density) and can be used in arbitrary volumetric renderers. Combining both representations (kernel-based representations for both media and radiance) would be an interesting line of future work.
@@ 3. Background
Under the geometric optics assumption, light transport in participating media is governed by the radiative transfer equation (RTE) (chandrasekhar1960radiative). We compute the incident radiance $L(\mathbf{x},\omega)$ at point $\mathbf{x}$ and direction $\omega$ as
@@
that models the light scattered in direction $\omega$ as a function of the integral of the light incoming from all directions on the unit sphere $\mathcal{S}^{2}$}. The interaction of light and matter is characterized by the optical properties of the medium, described in terms of the differential probabilities of absorption and scattering (or coefficients) $\mu_{a}(\mathbf{x})$ and $\mu_{s}(\mathbf{x})$ respectively, the extinction coefficient $\mu_{t}(\mathbf{x})=\mu_{a}(\mathbf{x})+\mu_{s}(\mathbf{x})$, the phase function $f_{\text{p}}(\mathbf{x},\omega^{\prime}\rightarrow\omega)$, and the emission $Q(\mathbf{x},\omega)$. Assuming uncorrelated statistics of matter, the differential probability of extinction is defined as the product between the density $\rho(\mathbf{x})$ of the particles forming the medium at point $\mathbf{x}$ [m-3] and the particles’ cross section $\sigma(\mathbf{x})$ [m2] as $\mu_{t}(\mathbf{x})=\rho(\mathbf{x})\,\sigma(\mathbf{x})$. For simplicity, we ignore the wavelength dependency in Equation (1), assume scalar elastic light-matter interactions, and omit the potential anisotropy of the medium (jakob2010radiative); generalizing the following derivations to include these additional dimensions is trivial. By integrating both sides of the differential RTE (1) along the direction $\omega$ we obtain the volumetric rendering equation, that models radiance $L(\mathbf{x},\omega)$ as
$L(\mathbf{x}_{0},\omega)$ $=\int_{0}^{s}\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})\,\mu_{t}(\mathbf{x}_{t})\,L_{o}(\mathbf{x}_{t},\omega)\,\text{d}t,$
(3) $+\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{s})L(\mathbf{x}_{s},\omega)\qquad\text{ with }$
@@
Finally, $L(\mathbf{x}_{s},\omega)$ is the outgoing radiance at the medium boundary $\mathbf{x}_{s}$ at distance $s$ modeled by the rendering equation (kajiya86).